How much pure acid should be mixed with four gallons of a 50% acid solution in order to get a 60% acid solution?
Answer by oberobic(2304)

With mixture problems, always keep track of how much 'pure' stuff you need.
In this case you have 50% solution, so it is 1/2 pure stuff and 1/2 other stuff.
You have 4 gallons, so the include 2 gallons of pure stuff and 2 gallons of other stuff (perhaps water).
.
You need 60% mixture.
And you have pure stuff to add.
.
So,
.
x = amount of pure stuff to add
.
50%(4) + 100%(x) = 60%(4+x)
.
Notice the final volume is 4+x.
.
.5(4) + 1(x) = .6(4) + .6(x)
.
multiply by 10 to eliminate decimals
.
5(4) + 10x = 6(4) + 6(x)
.
20 + 10x = 24 + 6x
.
4x = 4
.
x = 1
.
This suggests you add 1 gallon of pure stuff to 4 gallons of 50% mixture to end up with 5 gallons of 60% mixture.
.
Always check your solution.
.
5 * 60% = 3 gallons of pure stuff and 2 gallons of water
.
4 * 50% = 2 gallons of pure stuff
1 * 100% = 1 gallon of pure stuff
.
Correct!
.
Answer: Add 1 gallon of pure stuff.
.
Done.